# THIS CONTENT IS PROTECTED BY A PATENT

### v1

@@ 2.1 Foundations of the method
This form of the Boltzmann equation, although uncommon, is not new, it was previously discussed in the literature, for example by Wienke (1974, 1982). The well-known algorithm for x-rays, Acuros®(Vassiliev et al, 2010), also solves the Boltzmann equation, but in a different form that can be classified as Eulerian. Developers of Monte Carlo algorithms usually simply imitate the actual physical processes. However, Monte Carlo algorithms for particle transport can be formally derived as solvers of the integral form of the Boltzmann equation (Vassiliev, 2017).
Equation (1) is integro-differential. To solve it we use an iterative procedure based on a line integration method. We integrate Eq. (1) along the particle path step-by-step, starting at $t=0$ and making finite steps $\Delta t$ until all particles stop. The step size $\Delta t$ is not random. It is variable and optimized to achieve the best balance between computing speed and accuracy. The physics of proton interactions with matter is such that the difference between the proton path length $t$ and the corresponding penetration depth in matter $z$ is small. This difference is characterized by the detour factor defined as the ratio of projected range (i.e. average value of the depth to which a charged particle penetrates in the course of slowing down to rest) and continuous slowing down range (the total path length). For protons with energies from 10 MeV to 250 MeV propagating in water the detour factor is 0.9980-0.9989 (NIST 2022). That is, the difference between $t$ and $z$ is only 0.1-0.2%. Hence, we assume $t=z$. This approximation is applicable to other materials. For example, for protons with 1 mm range $z/t=0.998$ for ICRU compact boone, and $z/t=0.981$ for tungsten. The 2% difference fir tungsten can be corrected for by pretabulating $t(z)$ function. This approximation does not mean that we neglect lateral scattering of protons or recoil particles. Our solver calculates angular and radial distributions of proton fluence. Both distributions are very narrow. The approximation $t=z$ is even more accurate for heavier particles used in radiotherapy, such as helium and carbon ions. Therefore our methods can be extended to include such particles.
2.2 Multiple Coulomb scattering
2.2.1 Fluence spectra
@@
To calculate $\Phi_{i}(E|E^{\prime})$, we use Vavilov distribution (Vavilov 1957, Vassiliev 2017). Our algorithm allows to use Vavilov distribution for all steps, or only for a few first steps. Because calculation of Vavilov distribution is relatively slow, for optimal performance we recommend using it only for the first spectrum $\Phi_{1}(E|E^{\prime})$. Vavilov distribution has been previously used in advanced Monte Carlo software (Pelowitz 2011). Here we introduce a method for using it in a deterministic solver.
For all the steps that do not use Vavilov distribution, we approximate the conditional distribution $\Phi_{i}(E|E^{\prime})$ with a normal distribution. This approximation is based on the asymptotic properties of Vavilov distribution that tends to a normal distribution as the step size $\Delta t_{i}$ increases (Vavilov 1957, Vassiliev 2017). Hence, the step size $\Delta t_{i}$ should be sufficiently large for the normal distribution to be an accurate approximation of the exact Vavilov formula. On the other hand, calculations are substantially simplified when particle energy loss per step is small, which is achieved by limiting the step size from above. Balancing these two conflicting requirements determines the optimal step size. We determined optimal step sizes by performing our calculations with various step sizes and comparing the results with Monte Carlo simulations. Equation (1) does not account for loss of protons in nuclear reactions. We discuss this later.
Vavilov distribution. This is an overview only. For derivation of the distribution we refer to Vavilov (1957) and Vassiliev (2017). Vavilov distribution is a solution of the multiple scattering problem for charged particles that travel a distance $t$ such that the energy losses are much smaller than the initial energy of the particles. The solution is based on a relativistic form of the Rutherford formula for the scattering cross section (Vavilov 1957, Seltzer 1988):
@@
Normal distribution. If we approximate $\Phi_{i}(E|E^{\prime})$ with a normal distribution, then we need to calculate its center $\bar{E}$ and width $\sigma^{2}$. We use the continuous slowing down approximation (CSDA) to calculate $\bar{E}$. First, we tabulate range versus proton energy, $R(E)$. If a proton starts with an energy $E^{\prime}$ and travels distance $\Delta t$, then its average energy $\bar{E}$ at the step end is (Vassiliev, 2017):
$\bar{E}=\cases{R^{-1}\left[R\left(E^{\prime}\right)-\Delta t\right];&$\Delta t<R\left(E^{\prime}\right)$\\ 0;&$\Delta t\geq R\left(E^{\prime}\right),$\\ }$ (11)
where $R^{-1}$ is the inverse function. The corresponding distribution width is (Rotondi and Montagna 1990):
@@
2.2.2 Angular and radial distributions.
Iterative procedure for angular distributions. Angular variables can be separated from energy, because Cooulomb scattering is an elastic process. Angular distributions of fluence $\Phi(\vec{\Omega})$ are calculated iteratively using the following formula:
@@
In Eq. (21) we used a small angle approximation for $\theta^{\prime}$ (Vassiliev 2017). To calculate the integral over $\theta^{\prime}$ we use the Legendre quadrature. For the integral over $\phi^{\prime}$ we use the identity
@@
Molière distribution for a small step. To calculate angular distribution of protons for a small step $\Delta t$ we use Molière distribution. Here we give only a summary. Full details are given in Berger (1993) and Vassiliev (2017). The angular distribution is calculated as follows:
@@
Notations: $Z_{j}$ is the atomic number of element $j$; $\gamma$ is Euler’s constant; $\alpha$ is the fine structure constant; $C_{TF}=$0.88534 is the Thomas-Fermi constant; $k_{HF,j}$ is the Hartee-Fock correction for element $j$ (Berger 1993).
From angular distribution to radial distribution. Radial distribution of fluence $\Phi(z,\rho)$ is derived from angular distribution $\Phi(z,\vec{\Omega})$ using an adaptation for step-wise application of a technique introduced by Molière (1955) and discussed by Berger (1993). At depth $z_{0}=0$ we have $\bar{\rho}_{0}=\bar{\theta}_{0}=0$, where the bar indicates the average value. At depth $z_{1}>z_{0}$, we have $\bar{\rho}_{1}=\frac{1}{2}\tan{(\bar{\theta}_{1})}(z_{1}-z_{0})$. At depth $z_{i+1}$ we have:
@@
2.3 Nuclear interactions
The equation for secondary protons is a separate e quation, in which primary fluence is the source. Hence, nuclear processes are modelled after multiple Coulomb scattering calculations using Eq. (2) are completed. At that point we have proton fluence spectra $\Phi_{i}(E)$ calculated at all depths $z_{1},z_{2},\ldots,z_{n}$. We account for three nuclear processes: elastic interactions of protons with hydrogen atoms; elastic interactions of protons with atoms heavier than hydrogen (C,N,O, etc.); inelastic interactions of protons with atoms heavier than hydrogen. Inelastic interactions of protons with hydrogen atoms are negligible. The cross sections for these reactions are $\sigma_{el}^{H}(E)$, $\sigma_{el}^{A}(E)$, and $\sigma_{in}^{A}(E)$, respectively. The cross-sectional data were compiled from several sources: Barashenkov (1989), Arndt et al(1997, 1998), and Geant4 Collaboration (2018).
@@ 2.3 Nuclear interactions
2.3.1 Elastic scattering of protons on hydrogen atoms ($p\rightarrow H$)
@@
To calculate contributions to fluence spectra at a depth $\zeta$ from primary and recoil protons, $\Phi_{el,i}^{H}(\zeta,E)$, $i=1,2$, we use a method similar to the Monte Carlo surface tally (Vassiliev, 2017). First, we calculate distance $l_{i}$ in the direction defined by the directional cosine $\mu_{lab,i}$, from a given point in particle path, $t$, to the plane normal to the $z$-axis and located at depth $\zeta$:
@@
2.3.2 Elastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
Elastic scattering on atoms heavier than hydrogen is overall similar to $p\to H$ scattering. The main difference is that now we have one scattered proton instead of two. We do not consider displacement of a recoil atom $A$, and scattering is highly anisotropic. The differential cross section for targets with an atomic mass number $A<62$ is (Ranft 1972, Tseung et al2015):
@@ 2.3.2 Elastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
The total energy and momentum are transformed from the center of mass frame to the laboratory frame as follows (Geant4 1994):
@@
2.3.3 Inelastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
Cross sections for inelastic $p\to H$ interactions in tissue are zero for proton energies below 300 MeV. For heavier target atoms, after an inelastic $p\to A$ collision, the nucleus may emit gamma radiation, neutrons, low energy protons and heavy recoils (ICRU 2000). We account for energy transport only by protons. Heavy recoils are assumed to deposit energy locally, and neutral particles escape the volume without interactions. Total and single differential cross sections $\sigma_{in}^{A}(E)$ and $\sigma_{in}^{A}(E,\epsilon)$, respectively were taken from ICRU Report 63 (ICRU 2000). Contributions to proton fluence from nonelastic interactions were calculated by a formula similar to Eq. (54):
@@ 2.3.3 Inelastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
In this process the scattering angle does not uniquely determine the energy of the emitted proton $\epsilon$. For this reason Eq. (63) includes integration over $\epsilon$. In this equation we assumed isotropic angular distribution of emitted protons. A more accurate modelling of the angular distribution can be implemented using double differential cross sections that are also included in ICRU Report 63 (ICRU 2000). After preliminary testing we concluded that a good balance of computing speed and accuracy is achieved without modeling this process. It was not included in the calculations presented in this paper.
2.4 Algorithm implementation
2.4.1 Calculations in heterogeneous media
@@
2.4.2 Calculation of absorbed dose and average LET
@@ 2.4.5 Angular discretization
2.4.6 Discretization of radial distance
@@
3 Results and Discussion
The software is written in Fortran 95. We performed all the calculations on an HP Workstation with an Intel Core i7-10700 CPU, 2.9 GHz. For comparison, we repeated the same calculations using Monte Carlo software Geant4 with the physics list QGSP_BIC, optimal for hadron therapy (Geant4 Collaboration 2018). The medium in all the calculations is liquid water. Protons are incident normally on water surface.
3.1 Fluence spectra
We report results of fluence spectra calculations at various depths for a point monoenergetic monodirectional proton source ($\delta$-source). Energies of incident protons were 40 MeV (Fig. 3), 100 MeV (Fig. 4), 160 MeV (Fig. 5) and 220 MeV (Figs. 6). The CPU time for our DBS was 5-11 ms. CPU times for the Monte Carlo simulations were tens of hours. All the spectra are normalized per one incident proton. We made no other normalizations, scaling or adjustments of any kind.
@@ 3.2 Depth dose
3.3 Narrow Gaussian beam. Dose distributions
@@ 5 Acknowledgements
6 Competing Interests statement

### v2

@@ II.A. Foundations of the method
This form of the Boltzmann equation, although uncommon, is not new, it was previously discussed in the literature, for example by Wienke 42, 5. The well-known algorithm for x-rays, Acuros®39, also solves the Boltzmann equation, but in a different form that can be classified as Eulerian. Developers of Monte Carlo algorithms usually simply imitate the actual physical processes. However, Monte Carlo algorithms for particle transport can be formally derived as solvers of the integral form of the Boltzmann equation 40.
Equation (1) is integro-differential. To solve it we use an iterative procedure based on a line integration method. We integrate Eq. (1) along the particle path step-by-step, starting at $t=0$ and making finite steps $\Delta t$ until all particles stop. The step size $\Delta t$ is not random. It is variable and optimized to achieve the best balance between computing speed and accuracy. The physics of proton interactions with matter is such that the difference between the proton path length $t$ and the corresponding penetration depth in matter $z$ is small. This difference is characterized by the detour factor defined as the ratio of projected range (i.e. average value of the depth to which a charged particle penetrates in the course of slowing down to rest) and continuous slowing down range (the total path length). For protons with energies from 10 MeV to 250 MeV propagating in water the detour factor is 0.9980-0.9989 22. That is, the difference between $t$ and $z$ is only 0.1-0.2%. Hence, we assume $t=z$. This approximation is applicable to other materials. For example, for protons with 1 mm range $z/t=0.998$ for ICRU compact boone, and $z/t=0.981$ for tungsten. The 2% difference for tungsten can be corrected for by pretabulating $t(z)$ function. This approximation does not mean that we neglect lateral scattering of protons or recoil particles. Our solver calculates angular and radial distributions of proton fluence. Both distributions are very narrow. The approximation $t=z$ is even more accurate for heavier particles used in radiotherapy, such as helium and carbon ions. Therefore our methods can be extended to include such particles.
II.B. Multiple Coulomb scattering
II.B.1. Fluence spectra
@@
To calculate $\Phi_{i}(E|E^{\prime})$, we use Vavilov distribution 41, 40. Our algorithm allows to use Vavilov distribution for all steps, or only for a few first steps. Because calculation of Vavilov distribution is relatively slow, for optimal performance we recommend using it only for the first spectrum $\Phi_{1}(E|E^{\prime})$. Vavilov distribution has been previously used in advanced Monte Carlo software 25. Here we introduce a method for using it in a deterministic solver.
For all the steps that do not use Vavilov distribution, we approximate the conditional distribution $\Phi_{i}(E|E^{\prime})$ with a normal distribution. This approximation is based on the asymptotic properties of Vavilov distribution that tends to a normal distribution as the step size $\Delta t_{i}$ increases 41, 40. Hence, the step size $\Delta t_{i}$ should be sufficiently large for the normal distribution to be an accurate approximation of the exact Vavilov formula. On the other hand, calculations are substantially simplified when particle energy loss per step is small, which is achieved by limiting the step size from above. Balancing these two conflicting requirements determines the optimal step size. We determined optimal step sizes by performing our calculations with various step sizes and comparing the results with Monte Carlo simulations. Equation (1) does not account for loss of protons in nuclear reactions. We discuss this later.
Vavilov distribution. This is an overview only. For derivation of the distribution we refer to Vavilov41 and Vassiliev40. Vavilov distribution is a solution of the multiple scattering problem for charged particles that travel a distance $t$ such that the energy losses are much smaller than the initial energy of the particles. The solution is based on a relativistic form of the Rutherford formula for the scattering cross section 41, 35:
@@
Normal distribution. If we approximate $\Phi_{i}(E|E^{\prime})$ with a normal distribution, then we need to calculate its center $\bar{E}$ and width $\sigma^{2}$. We use the continuous slowing down approximation (CSDA) to calculate $\bar{E}$. First, we tabulate range versus proton energy, $R(E)$. If a proton starts with an energy $E^{\prime}$ and travels distance $\Delta t$, then its average energy $\bar{E}$ at the step end is40:
$\bar{E}=\begin{cases}R^{-1}\left[R\left(E^{\prime}\right)-\Delta t\right];&\Delta t<R\left(E^{\prime}\right)\\ 0;&\Delta t\geq R\left(E^{\prime}\right),\end{cases}$
where $R^{-1}$ is the inverse function. The corresponding distribution width is29 :
@@
II.B.2. Angular and radial distributions.
Iterative procedure for angular distributions. Angular distributions of fluence $\Phi(\vec{\Omega})$ are calculated iteratively using the following formula:
@@
In Eq. (21) we used a small angle approximation for $\theta^{\prime}$ 40. To calculate the integral over $\theta^{\prime}$ we use the Legendre quadrature. For the integral over $\phi^{\prime}$ we use the identity
@@
Molière distribution for a small step. To calculate angular distribution of protons for a small step $\Delta t$ we use Molière distribution. Here we give only a summary. Full details are given in Berger8 and Vassiliev40. The angular distribution is calculated as follows:
@@
Notations: $Z_{j}$ is the atomic number of element $j$; $\gamma$ is Euler’s constant; $\alpha$ is the fine structure constant; $C_{TF}=$0.88534 is the Thomas-Fermi constant; $k_{HF,j}$ is the Hartee-Fock correction for element jj8.
From angular distribution to radial distribution. Radial distribution of fluence $\Phi(z,\rho)$ is derived from angular distribution $\Phi(z,\vec{\Omega})$ using an adaptation for step-wise application of a technique introduced by Molière 20 and discussed by Berger8. At depth $z_{0}=0$ we have $\bar{\rho}_{0}=\bar{\theta}_{0}=0$, where the bar indicates the average value. At depth $z_{1}>z_{0}$, we have $\bar{\rho}_{1}=\frac{1}{2}\tan{(\bar{\theta}_{1})}(z_{1}-z_{0})$. At depth $z_{i+1}$ we have:
@@
II.C. Nuclear interactions
The equation for secondary protons is a separate e quation, in which primary fluence is the source. Hence, nuclear processes are modelled after multiple Coulomb scattering calculations using Eq. (2) are completed. At that point we have proton fluence spectra $\Phi_{i}(E)$ calculated at all depths $z_{1},z_{2},\ldots,z_{n}$. We account for three nuclear processes: elastic interactions of protons with hydrogen atoms; elastic interactions of protons with atoms heavier than hydrogen (C,N,O, etc.); inelastic interactions of protons with atoms heavier than hydrogen. Inelastic interactions of protons with hydrogen atoms are negligible. The cross sections for these reactions are $\sigma_{el}^{H}(E)$, $\sigma_{el}^{A}(E)$, and $\sigma_{in}^{A}(E)$, respectively. The cross-sectional data were compiled from several sources V. S. Barashenkov (1989), R. A. Arndt et al. (1997), R. A. Arndt (1998), 13.
@@ II.C. Nuclear interactions
II.C.1. Elastic scattering of protons on hydrogen atoms ($p\rightarrow H$)
Nuclear reactions contribute much less to dose than do Coulomb interactions with atomic electrons. Therefore, we use a different, more sparse, grid $\zeta_{1},\ldots,\zeta_{n}$ for nuclear reactions than $z_{1},\ldots,z_{n}$ that we use for multiple Coulomb scattering. How these grids are designed is discussed in Subsection II.D.3.
Angular distributions of scattered protons in the center of mass frame are shown in Fig. 2. They are taken from Arndt4. We use the isotropic scattering approximation $f_{0}(\mu_{cm})=1/2$, where $\mu_{cm}$ is the cosine of scattering angle in the center of mass frame and $f_{0}(\mu_{cm})$ is the probability density. The latter is shown with a red dashed line in Fig. 2. This is a sufficiently accurate approximation, because this process is relatively rare. If a more accurate model of this process is needed, a weight $w=f(\mu_{cm})/f_{0}(\mu_{cm})$ is assigned to the particle when it scatters by an angle $\mu_{cm}$, where $f(\mu_{cm})$ is the accurate angular distribution (Fig. 2).
@@ II.C.1. Elastic scattering of protons on hydrogen atoms ($p\rightarrow H$)
To calculate contributions to fluence spectra at a depth $\zeta$ from primary and recoil protons, $\Phi_{el,i}^{H}(\zeta,E)$, $i=1,2$, we use a method similar to the Monte Carlo surface tally40. First, we calculate distance $l_{i}$ in the direction defined by the directional cosine $\mu_{lab,i}$, from a given point in particle path, $t$, to the plane normal to the $z$-axis and located at depth $\zeta$:
@@
II.C.2. Elastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
Elastic scattering on atoms heavier than hydrogen is overall similar to $p\to H$ scattering. The main difference is that now we have one scattered proton instead of two. We do not consider displacement of a recoil atom $A$, and scattering is highly anisotropic. The differential cross section for targets with an atomic mass number $A<62$ is 27, 37:
@@ II.C.2. Elastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
The total energy and momentum are transformed from the center of mass frame to the laboratory frame as follows14:
@@
II.C.3. Inelastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
Cross sections for inelastic $p\to H$ interactions in tissue are zero for proton energies below 300 MeV. For heavier target atoms, after an inelastic $p\to A$ collision, the nucleus may emit gamma radiation, neutrons, low energy protons and heavy recoils16. We account for energy transport only by protons. Heavy recoils are assumed to deposit energy locally, and neutral particles escape the volume without interactions. Total and single differential cross sections $\sigma_{in}^{A}(E)$ and $\sigma_{in}^{A}(E,\epsilon)$, respectively were taken from ICRU Report 6316. Contributions to proton fluence from nonelastic interactions were calculated by a formula similar to Eq. (54):
@@ II.C.3. Inelastic scattering of protons on atoms heavier than hydrogen ($p\rightarrow A$)
In this process the scattering angle does not uniquely determine the energy of the emitted proton $\epsilon$. For this reason Eq. (63) includes integration over $\epsilon$. In this equation we assumed isotropic angular distribution of emitted protons. A more accurate modelling of the angular distribution can be implemented using double differential cross sections that are also included in ICRU Report 63 16. After preliminary testing we concluded that a good balance of computing speed and accuracy is achieved without modeling this process. It was not included in the calculations presented in this paper.
II.D. Algorithm implementation
II.D.1. Calculations in heterogeneous media
@@
II.D.2. Calculation of absorbed dose and average LET
@@ II.D.5. Angular discretization
II.D.6. Discretization of radial distance
@@
III. Results and Discussion
The software is written in Fortran 95. We performed all the calculations on an HP Workstation with an Intel Core i7-10700 CPU, 2.9 GHz. For comparison, we repeated the same calculations using Monte Carlo software Geant4 with the physics list QGSP_BIC, optimal for hadron therapy13. The medium in all the calculations is liquid water. Protons are incident normally on water surface.
III.A. Fluence spectra
We report results of fluence spectra calculations at various depths for a point monoenergetic monodirectional proton source ($\delta$-source). Energies of incident protons were 40 MeV (Fig. 3), 100 MeV (Fig. 4), 160 MeV (Fig. 5) and 220 MeV (Fig. 6). The CPU time for our DBS was 5-11 ms. CPU times for the Monte Carlo simulations were tens of hours. All the spectra are normalized per one incident proton. We made no other normalizations, scaling or adjustments of any kind.
@@ III.B. Depth dose
III.C. Narrow Gaussian beam. Dose distributions
@@ V. Acknowledgements
VI. Competing Interests statement